# DMesh: A Differentiable Mesh Representation

### v2

@@ Abstract
We present a differentiable representation, DMesh, for general 3D triangular meshes. DMesh considers both the geometry and connectivity information of a mesh. In our design, we first get a set of convex tetrahedra that compactly tessellates the domain based on Weighted Delaunay Triangulation (WDT), and select triangular faces on the tetrahedra to define the final mesh. We formulate probability of faces to exist on the actual surface in a differentiable manner based on the WDT. This enables DMesh to represent meshes of various topology in a differentiable way, and allows us to reconstruct the mesh under various observations, such as point cloud and multi-view images using gradient-based optimization. The source code and full paper is available at: https://sonsang.github.io/dmesh-project  [^1] This paper was last modified at May 31, 2024.
@@ 1 Introduction
  • We propose a computationally efficient approach to differentiable WDT, which produces robust probability estimations. While exhaustive approach (Rakotosaona u. a., 2021) requires exponential computational cost, our method runs in approximately linear time.
@@ 2.2 Delaunay Triangulation for Geometry Processing
More recently, (Rakotosaona u. a., 2021) focused on this DT’s property to connect points and tessellate the domain, and proposed a differentiable WDT algorithm to compute smooth inclusion, namely existence score of $2$-simplexes (triangles) in 2 dimensional space. However, it is not suitable to apply this approach to our 3D case, as there are computational challenges (Section 3.2). Other related work, VoroMesh (Maruani u. a., 2023), also used Voronoi diagrams in point cloud reconstruction, but their formulation cannot represent open surfaces and is only confined to handle point clouds.
@@ 3.1 Probabilistic Approach to Mesh Connectivity
$\Lambda(F)=P(F\in Mesh)=\Lambda_{wdt}(F)\cdot\Lambda_{real}(F).$ (2)
This formulation attains one nice property in determining the final mesh – that is, it prohibits self-intersections between faces. When it comes to the other two criteria about oracle, $\Lambda_{wdt}$ function’s differentiability and efficiency is crucial, as we design $\Lambda_{real}$ to be a very efficient differentiable function based on real values (Section 4.2). Thus, we first introduce how we can evaluate $\Lambda_{wdt}$ in a differentiable and efficient manner, which is one of our main contributions.
@@ 3.2 Basic Principles
To begin with, we use $(d,k)$ pair to denote a $k$-simplex ($\Delta^{k}$) in $d$-dimensional space. For a 3D mesh, we observe that our face $F$ corresponds to $(d=3,k=2)$ in Figure 5(b). To compute the probability $\Lambda_{wdt}$ for the face, we use power diagram (PD), which is the dual structure of WDT (Figure 6(a)). While previously Rakotosaona et al. (2021) proposed a differentiable 2D triangulation method for the $(d=k=2)$ case (Figure 5(c)), it suffers from exponentially increasing computational cost (e.g. it takes 4.3 seconds to process $10K$ points in 2D, Table 4) and unreliable estimation when $(k<d)$. We will discuss later how our formulation conquer these computational challenges. Our setting for 3D meshes are similar to 2D meshes, the $(d=2,k=1)$ case in Figure 5(d), where a triangular face reduces to a line. Therefore, we will mainly use this setting to describe and visualize basic concepts for simplicity. However, note that it can be generalized to any $(d,k)$ case without much problem. We generalize the basic principles suggested by Rakotosaona u. a. (2021) to address our cases. For formal definitions of the concepts in this section, please refer to Appendix B.
@@
To make this measurement less binary when $\Delta^{1}$ exists, we use the expanded version of power cell called "reduced" power cell ($R_{p|\Delta}$), introduced by Rakotosaona u. a. (2021). The reduced power cell of $p_{i}\in S$ for $\Delta^{1}=\{p_{i},p_{j}\}$ is computed by excluding $p_{j}$ from $S$ when constructing the power cell  [^4] In 3D case where $k=2$, $\Delta^{2}=\{p_{i},p_{j},p_{k}\}$, and $p_{j}$ and $p_{k}$ will be ignored for $R_{p_{i}|\Delta^{2}}$.. For example, when $\Delta^{1}$ exists, $R_{p_{i}|\Delta^{1}}$ will expand towards $p_{j}$’s direction (Figure 6(c)), and $D_{\Delta^{1}}$ will "go through" $R_{p_{i}|\Delta^{1}}$. In contrast, when $\Delta$ doesn’t exists, even though we have removed $p_{j}$, $R_{p_{i}|\Delta^{1}}$ will not expand (Figure 6(d)), and thus $D_{\Delta^{1}}$ stays outside of $R_{p_{i}|\Delta^{1}}$.
@@
  • Efficiency: When naïvely estimating the reduced power cell in exhaustive manner, the computational cost increases exponentially with the number of points at a rate of $O(N^{2})$, where $N$ is the number of points.
@@ 4.2 Definition of $\Lambda_{real}$
$\Lambda_{real}$ evaluates the existence probability of a $k$-simplex $\Delta^{k}$ in our mesh when it exists in WDT. To define it, we leverage per-point value $\psi\in[0,1]$. We simply use the minimum $\psi$ of the points in $\Delta^{k}$: $\Lambda_{real}(\Delta^{k})=\min_{p\in\Delta^{k}}\Psi(p)$, where $\Psi$ is function that maps a point $p$ to its $\psi$ value.
@@ 4.3.2 Regularizations
Being fully differentiable for both vertices and faces, DMesh allows us to develop various regularizations to improve the optimization process and enhance the final mesh quality. The first is weight regularization ($L_{weight}$), applied to the dual Power Diagram of the WDT (Appendix C.4). This regularization reduces the structural complexity of the WDT and discards unnecessary points, controlling the final mesh complexity (Figure 7). The next is real regularization ($L_{real}$), which enforces nearby points to have similar real values and increases the real values of points adjacent to high real value points (Appendix C.5). This helps remove holes or inner structures and makes faces near the current surface more likely to be considered (Appendix D). The final regularization, quality regularization ($L_{qual}$), aims to improve the quality of triangle faces by minimizing the average expected aspect ratio of the faces, thus removing thin triangles (Appendix C.6).
@@ 5.2 Point Cloud & Multi-View Reconstruction
6 Conclusion and Future Directions
Our method achieves a more effective and complete forms of differentiable meshes of various topology than existing methods, but opens up new avenues for future research.
  • Computational cost: Current DMesh resolution is constrained by computational cost. Although we succeeded in decreasing computational burden through our theoretical relaxation and CUDA implementation, it takes more than a second to process over 100K vertices in 3D due to costly WDT construction.
  • Non-manifoldness: As we have claimed so far, DMesh shows much better generalization than the other methods as it does not have any constraints on the mesh connectivity. However, due to this relaxation of constraint, we can observe spurious non-manifold errors in the mesh, even though we adopted measures to minimize them (Appendix D.2.7).
To address these limitations, we can search for a method that does not heavily rely on WDT algorithm as we do now, as it requires more computational cost when the number of points exceeds the on-chip memory of GPU support. We may also adopt the other approaches based on (un)signed distance fields (Shen u. a., 2023; Liu u. a., 2023b) to our method, thus generating compact mesh that is guaranteed to be manifold. Finally, further research can extend this work to solve other challenging problems (e.g. 3D reconstruction from real-world images) or other related applications (e.g. 3D mesh generative model) in the future.
@@ B.1 Mathematical Definitions
Here, we provide formal mathematical definitions of the terms used in Section 3.2. We mainly use notations from Aurenhammer u. a. (2013); Cheng u. a. (2013).
@@ E.2.3 Visual Renderings
In Figure 2021, and 22, we provide visual renderings of our point cloud and multi-view reconstruction results with ground truth mesh. We also provide illustration of input point cloud and diffuse map. Note that we also used depth renderings for multi-view reconstruction experiments.

### v3

@@ Abstract
We present a differentiable representation, DMesh, for general 3D triangular meshes. DMesh considers both the geometry and connectivity information of a mesh. In our design, we first get a set of convex tetrahedra that compactly tessellates the domain based on Weighted Delaunay Triangulation (WDT), and select triangular faces on the tetrahedra to define the final mesh. We formulate probability of faces to exist on the actual surface in a differentiable manner based on the WDT. This enables DMesh to represent meshes of various topology in a differentiable way, and allows us to reconstruct the mesh under various observations, such as point clouds and multi-view images using gradient-based optimization. We publicize the source code and supplementary material at our project page  [^1] https://sonsang.github.io/dmesh-project.
@@ 1 Introduction
  • We propose a computationally efficient approach to differentiable WDT, which produces robust probability estimations. While exhaustive approach (Rakotosaona u. a., 2021) requires quadratic computational cost, our method runs in approximately linear time.
@@ 2.2 Delaunay Triangulation for Geometry Processing
More recently, Rakotosaona u. a. (2021) focused on this DT’s property to connect points and tessellate the domain, and proposed a differentiable WDT algorithm to compute smooth inclusion, namely existence score of $2$-simplexes (triangles) in 2 dimensional space. However, it is not suitable to apply this approach to our 3D case, as there are computational challenges (Section 3.2). Other related work, VoroMesh (Maruani u. a., 2023), also used Voronoi diagrams in point cloud reconstruction, but their formulation cannot represent open surfaces and is only confined to handle point clouds.
@@ 3.1 Probabilistic Approach to Mesh Connectivity
$\Lambda(F)=P(F\in\text{Mesh})=\Lambda_{wdt}(F)\cdot\Lambda_{real}(F).$ (2)
This formulation attains one nice property in determining the final mesh – that is, it prohibits self-intersections between faces. When it comes to the other two criteria about this procedure, $\Lambda_{wdt}$ function’s differentiability and efficiency is crucial, as we design $\Lambda_{real}$ to be a very efficient differentiable function based on real values (Section 4.2). Thus, we first introduce how we can evaluate $\Lambda_{wdt}$ in a differentiable and efficient manner, which is one of our main contributions.
@@ 3.2 Basic Principles
To begin with, we use $(d,k)$ pair to denote a $k$-simplex ($\Delta^{k}$) in $d$-dimensional space. For a 3D mesh, we observe that our face $F$ corresponds to $(d=3,k=2)$ in Figure 5(b). To compute the probability $\Lambda_{wdt}$ for the face, we use power diagram (PD), which is the dual structure of WDT (Figure 6(a)). While previously Rakotosaona et al. (2021) proposed a differentiable 2D triangulation method for the $(d=k=2)$ case (Figure 5(c)), it suffers from quadratically increasing computational cost (e.g. it takes 4.3 seconds to process $10K$ points in 2D, Table 4) and unreliable estimation when $(k<d)$. We will discuss later how our formulation conquer these computational challenges. Our setting for 3D meshes are similar to 2D meshes, the $(d=2,k=1)$ case in Figure 5(d), where a triangular face reduces to a line. Therefore, we will mainly use this setting to describe and visualize basic concepts for simplicity. However, note that it can be generalized to any $(d,k)$ case without much problem. We generalize the basic principles suggested by Rakotosaona u. a. (2021) to address our cases. For formal definitions of the concepts in this section, please refer to Appendix B.
@@
To make this measurement less binary when $\Delta^{1}$ exists, we use the expanded version of power cell called "reduced" power cell ($R_{p|\Delta}$), introduced by Rakotosaona u. a. (2021). The reduced power cell of $p_{i}\in S$ for $\Delta^{1}=\{p_{i},p_{j}\}$ is computed by excluding $p_{j}$ from $S$ when constructing the power cell  [^4] In 3D case where $k=2$ and $\Delta^{2}=\{p_{i},p_{j},p_{k}\}$, $p_{j}$ and $p_{k}$ would be ignored for $R_{p_{i}|\Delta^{2}}$.. For example, when $\Delta^{1}$ exists, $R_{p_{i}|\Delta^{1}}$ will expand towards $p_{j}$’s direction (Figure 6(c)), and $D_{\Delta^{1}}$ will "go through" $R_{p_{i}|\Delta^{1}}$. In contrast, when $\Delta$ doesn’t exists, even though we have removed $p_{j}$, $R_{p_{i}|\Delta^{1}}$ will not expand (Figure 6(d)), and thus $D_{\Delta^{1}}$ stays outside of $R_{p_{i}|\Delta^{1}}$.
@@
  • Efficiency: When naïvely estimating the reduced power cell in exhaustive manner, the computational cost increases with the number of points at a rate of $O(N^{2})$, where $N$ is the number of points.
@@ 4.2 Definition of $\Lambda_{real}$
$\Lambda_{real}$ evaluates the existence probability of a $k$-simplex $\Delta^{k}$ in our mesh when it exists in WDT. To define it, we leverage per-point value $\psi\in[0,1]$. To be specific, we compute the minimum $\psi$ of the points in $\Delta^{k}$ in differentiable way: $\Lambda_{real}(\Delta^{k})=\sum_{p\in\Delta^{k}}\kappa(p)\cdot\Psi(p)$, where $\kappa(p)=e^{-\beta\cdot\Psi(p)}/\sum_{q\ \in\Delta^{k}}e^{-\beta\cdot\Psi(q)}$, and $\Psi$ is function that maps a point $p$ to its $\psi$ value. We set $\beta=100$.
@@ 4.3.2 Regularizations
Being fully differentiable for both vertices and faces, DMesh allows us to develop various regularizations to improve the optimization process and enhance the final mesh quality. The first is weight regularization ($L_{weight}$), applied to the dual Power Diagram of the WDT (Appendix C.4). This regularization reduces the structural complexity of the WDT, controlling the final mesh complexity (Figure 7). The next is real regularization ($L_{real}$), which enforces nearby points to have similar real values and increases the real values of points adjacent to high real value points (Appendix C.5). This helps remove holes or inner structures and makes faces near the current surface more likely to be considered (Appendix D). The final regularization, quality regularization ($L_{qual}$), aims to improve the quality of triangle faces by minimizing the average expected aspect ratio of the faces, thus removing thin triangles (Appendix C.6).
@@ 5.2 Point Cloud & Multi-View Reconstruction
Figure 9: ()($\rightarrow$)( → ) Shape interpolation using DMesh exhibiting topology change. After fitting DMesh to a torus (upper left), we optimize it again to reconstruct a double torus (lower right), which has a different genus. We use multi-view images for the optimization.
@@
Before moving on, we present an experimental result about shape interpolation using DMesh in Figure 9. We used multi-view images to reconstruct a torus first, and then optimized the DMesh again to fit a double torus. The results show that DMesh effectively reconstructs the double torus, even when initialized from a converged single torus, highlighting the method’s robustness to local minima. However, this also indicates that our representation lacks meaningful shape interpolation, as it does not assume any specific shape topology.
6 Conclusion
6.1 Limitations
As shown above, our method achieves a more effective and complete forms of differentiable meshes of various topology than existing methods, but still has several limitations to overcome.
Figure 10: Analysis of computational cost for computing face existence probabilities ($\Lambda(F)$). The computational cost rises sharply beyond 20K points, with most of the time spent on WDT construction (“WDT"), while the probability computation (“Prob") requires significantly less time.
Computational Cost. Currently, the resolution of DMesh is limited by computational cost. Although our theoretical relaxation and CUDA implementation reduce this burden, processing meshes with over 100K vertices remains challenging due to the computational bottleneck of constructing the WDT at each optimization step. In Figure 10, we analyze computational costs relative to the number of points. As shown, costs rise sharply beyond 20K points, with WDT construction consuming most of the time. This limits our method’s ability to handle high resolution mesh.
Non-Manifoldness. As we have claimed so far, DMesh shows much better generalization than the other methods as it does not have any constraints on the shape topology and mesh connectivity. However, due to this relaxation of constraint, we can observe spurious non-manifold errors in the mesh, even though we adopted measures to minimize them (Appendix D.2.7).
Specifically, an edge must have at most two adjacent faces to be a "manifold" edge. Similarly, a "manifold" vertex should be adjacent to a set of faces that form a closed or open fan. We refer to edges or vertices that do not satisfy these definitions as "non-manifold." In our results, we found that 5.50% of edges and 0.38% of vertices were non-manifold for point cloud reconstruction. For multi-view reconstruction, 6.62% of edges and 0.25% of vertices were non-manifold. Therefore, we conclude that non-manifold edges are more prevalent than non-manifold vertices in our approach.
6.2 Future Work
To address the computational cost issue, we can explore methods that reduce reliance on the WDT algorithm, as its cost increases significantly with the number of points. This is crucial since representing complex shapes with fine details often requires over 100K vertices. To tackle the non-manifold issue, we could integrate approaches based on (un)signed distance fields (Shen u. a., 2023; Liu u. a., 2023b) into our method, ensuring manifold mesh generation. Finally, future research could extend this work to solve other challenging problems, such as 3D reconstruction from real-world images, or applications like generative models for 3D shapes. This could involve encoding color or texture information within our framework, opening up exciting new directions for exploration.
@@ B.1 Mathematical Definitions
Here, we provide formal mathematical definitions of the terms used in Section 3.2. We mainly use notations from Aurenhammer u. a. (2013) and  Cheng u. a. (2013).
@@ E.2.3 Visual Renderings
In Figure 2223, and 24, we provide visual renderings of our point cloud and multi-view reconstruction results with ground truth mesh. We also provide illustration of input point cloud and diffuse map. Note that we also used depth renderings for multi-view reconstruction experiments.